OpenStudy (anonymous):

What is the fifth term of the given sequence? 1/2x^2y, x^3, (2x^4)/y,....

OpenStudy (anonymous):

can you write your function here please

OpenStudy (anonymous):

There isn't a function????

OpenStudy (anonymous):

the way you've written your question its sorta confusing can u rewrite it please

OpenStudy (anonymous):

What is the fifth term of the geometric sequence when a1 = 1/2x^2y, a2 = x^3, a3 = (2x^4)/y

OpenStudy (anonymous):

\[a1=\frac{ 1 }{ 2x^{2y} }, a2=x^3,a3=\frac{ 2x^4 }{ y }\] is this wat u mean?

OpenStudy (anonymous):

1/2 * x^2y

OpenStudy (anonymous):

im sorry babe i aint getting your question

OpenStudy (anonymous):

Never mind then

OpenStudy (phi):

it looks like you multiply by \[ \frac{2x}{y} \] to get to the next term

OpenStudy (phi):

multiply the 3rd term by 2x/y and then multiply by 2x/y again to get the 5th term. that means multiply the 3rd term by \[ \frac{4x^2}{y^2} \] to get the answer

OpenStudy (anonymous):

Sooo... 8x^6 / y^3???

OpenStudy (phi):

I have not checked... \[ \frac{2x^4}{y} \cdot \frac{4x^2}{y^2} \] that looks like \[ \frac{8x^6}{y^3} \]

OpenStudy (anonymous):

I have done it right then! :D

OpenStudy (phi):

yes, looks good